# Many-Shot In-Context Learning

### v1

@@ Abstract
Large language models (LLMs) excel at few-shot in-context learning (ICL) – learning from a few examples provided in context at inference, without any weight updates. Newly expanded context windows allow us to investigate ICL with hundreds or thousands of examples – the many-shot regime. Going from few-shot to many-shot, we observe significant performance gains across a wide variety of generative and discriminative tasks. While promising, many-shot ICL can be bottlenecked by the available amount of human-generated outputs. To mitigate this limitation, we explore two new settings: “Reinforced ICL” and “Unsupervised ICL”. Reinforced ICL uses model-generated chain-of-thought rationales in place of human rationales. Unsupervised ICL removes rationales from the prompt altogether, and prompts the model only with domain-specific inputs. We find that both Reinforced and Unsupervised ICL can be quite effective in the many-shot regime, particularly on complex reasoning tasks. Finally, we demonstrate that, unlike few-shot learning, many-shot learning is effective at overriding pretraining biases and can learn high-dimensional functions with numerical inputs. Our analysis also reveals the limitations of next-token prediction loss as an indicator of downstream performance.
@@ 1 Introduction
Figure 1: Many-shot vs Few-Shot In-Context Learning (ICL) across several tasks. Many-shot learning exhibits consistent performance gains over few-shot ICL. This gain is especially dramatic for difficult non-natural language tasks like sequential parity prediction and linear classification. Number of best-performing shots for many-shot ICL are shown inside the bar for each task. For few-shot ICL, we either use typical number of shots used on a benchmark, for example, 4-shot for MATH, or the longest prompt among the ones we tested with less than the GPT-3 context length of 2048 tokens. Reasoning-oriented tasks, namely MATH, GSM8K, BBH, and GPQA uses human-generated chain-of-thought rationales. For translation, we report performance FLORES-MT result on English to Kurdish, summarization uses XLSum, MATH corresponds to the MATH500 test set, and sentiment analysis results are reported with semantically-unrelated labels. See §3, §4, and §5 for more details.
Large language models (LLMs) have demonstrated a remarkable ability to perform in-context learning (ICL): they can learn a new task just from input-output examples, also known as shots, which precede a test input presented within the LLM context. However, an LLM’s context window, i.e. the amount of tokenized inputs they can process for each forward inference, limits the number of shots that can be used for ICL, restricting prior research to the few-shot learning regime.
@@
While relatively under-explored, many-shot learning – ICL with a large number of shots, for example, hundreds or more – can potentially make LLMs more versatile, adaptable, and ultimately more useful. For example, many shots allow clearer specification of the task, whereas doing so through few shots may be more ambiguous. Exploring many-shot ICL is now feasible, given the recent increase in context windows of publicly available LLMs by at least $100\times$: from only a few thousand tokens in GPT-3 (Brown et al., 2020) and Llama 2 (Touvron et al., 2023) to 1M tokens in Gemini 1.5 Pro (Gemini Team, 2024).
@@
In this paper, we investigate how scaling the number of in-context examples (shots) effects LLM performance across diverse downstream tasks. To gain broad insights into many-shot learning, we analyze its effects on a variety of widely-used tasks: math problem solving using MATH (Hendrycks et al., 2021) and GSM8K (Cobbe et al., 2021), question-answering (GPQA, Rein et al., 2023), summarization using XSum (Narayan et al., 2018) and XLSum (Hasan et al., 2021), algorithmic reasoning (Big-Bench Hard, Suzgun et al., 2022), outcome reward modeling (Code verification, Ni et al., 2023), low-resource machine translation (FLORES, Goyal et al., 2022), planning (Logistics, Seipp et al., 2022), and sentiment analysis (FP, Malo et al., 2014). Compared to few-shot learning, many-shot learning leads to significant performance gains across these tasks, using several hundreds or thousands of shots with Gemini 1.5 Pro (Figure 1). Furthermore, maximum performance is often achieved only once the number of shots reaches up to hundreds of thousands of tokens (Figure 2).
@@
While many-shot learning holds significant promise, it can be constrained by the need for high-quality, human-generated outputs. This limitation is amplified in complex reasoning tasks, such as GPQA. To overcome this, we introduce reinforced ICL and unsupervised ICL. Inspired by the efficacy of model-generated solutions for fine-tuning (Singh et al., 2023), Reinforced ICL involves replacing human-written rationales with model-generated ones, filtered via answer correctness, for in-context learning. Inspired by task-recognition view of ICL (Xie et al., 2021), we also introduce Unsupervised ICL where we prompt the model with only problems instead of problem-solution pairs. On reasoning and problem-solving tasks, such as MATH and GPQA, we find that both reinforced and unsupervised ICL with many-shots can be more effective than few-shot ICL with human-generated rationales.
@@
Finally, we empirically study how the learning dynamics of in-context learning changes from few-shot to the many-shot regime. We find that with sufficient examples, ICL can overcome pre-training biases and solve high-dimensional prediction tasks with numerical inputs, namely sequential parity prediction and linear classification. This suggests the potential of many-shot learning to adapt to new tasks and domains that might be misaligned with an LLM’s training data. Surprisingly, the order of examples can significantly influence ICL performance, even in the many-shot setting. Furthermore, we discover that the widely-used next-token prediction loss may not reliably predict ICL performance on problem-solving and reasoning tasks.
@@
  • Scaling In-Context Learning (ICL)3): We systematically evaluate LLM performance at different scales of in-context examples for a wide range of tasks. Our results indicate significant performance leaps for LLMs transitioning from few-shot to many-shot learning regime.
  • Reinforced and Unsupervised ICL4): We find that either using model-generated rationales or only problems can reduce the dependence of many-shot learning on human-generated data.
  • Analysing ICL5): We find that many-shot ICL can overcome pre-training biases and learn non-natural language prediction tasks, where few-shot ICL struggles.
2 Related Work
Scaling in-context learning
Brown et al. (2020) reported improved performance as you increase the number of examples used for in-context learning in large language models, and later works corroborated this finding (Lu et al., 2022). However, very few works have explored using a large number of examples (1000 or above) in the prompt. This is likely due to the fact the context lengths in large language models have been quite limited until recently (Gemini Team, 2024; Anthropic, 2024). One closely related work to ours is from Li et al. (2023), who scale the number of examples for in-context learning to 2000. However, Li et al. (2023) use a custom model architecture (Zheng et al., 2023) to achieve long context lengths, and only evaluate models of up to 1.3B parameters, which is several orders of magnitude smaller than state-of-the-art language models, and are ineffective for complex tasks, such as GPQA (Rein et al., 2023).
Concurrently to our work, Anil et al. (2024) used many-shot prompting (upto 256 shots) to jailbreak language models. In our work, we focus on a much wider range of tasks, use a lot more examples (up to 8192 shots) and use models with much longer context lengths (up to 1M tokens). Also, we explore mitigations for needing many human-generated examples with many-shot ICL. Furthermore, while Anil et al. (2024) use many-shot learning to override preferences learned during RLHF phase to elicit the biases stemming from pretraining, our results in §5.1 demonstrate that we can also override pretraining biases themselves.
Long-context scaling laws
Prior works (Xiong et al., 2023; Anil et al., 2024; Kaplan et al., 2020; Gemini Team, 2024) have reported smaller next-token prediction loss with longer contexts, which Jeon et al. (2024) also show using theoretical analysis. Our findings confirm this trend for even longer context lengths, but our analysis reveals some of the limitations of using next-token prediction loss as a metric for evaluating long-context performance, as next-token prediction loss continues to go down even as overall performance plateaus.
Learning from self-generated data
Numerous recent works (Gulcehre et al., 2023; Yuan et al., 2023; Singh et al., 2023) propose fine-tuning language models on self-generated data to improve performance. Their approach consists of (1) generate samples from the model and filter them using binary feedback, (2) fine-tune the model on these samples, and (3) repeat this process a few times. In this work, we extend this idea to the domain of in-context learning, and study the efficacy of Reinforced ICL in reasoning and problem-solving domains.
Self-generated data and in-context learning
Kim et al. (2022) explore performing in-context learning using self-generated data for classification problems. Their approach consists of generating demonstrations conditioned on the test input for each possible class label, and including these demonstrations in the context when performing the final prediction. In contrast, our approach generates data using the same procedure as Singh et al. (2023), does not require access to the test input for data generation, and can be applied to any problem for which we can obtain reliable reward signals (not just classification problems).
Learning Input-Output Relationships with ICL
Numerous works (Min et al., 2022; Kossen et al., 2023; Yoo et al., 2022; Lin and Lee, 2024) have investigated whether LLMs truly learn input-output relationships during in-context learning. Min et al. (2022) found that replacing the ground truth labels in in-context examples with random labels barely effected final performance. Further investigations by Yoo et al. (2022) and Kossen et al. (2023) found that this finding does not necessarily hold across tasks and model sizes. In particular, Kossen et al. (2023); Lin and Lee (2024) showed that LLMs can indeed learn input-output relationships via in-context learning, but require more examples in order to do so well. In our work, we extrapolate the trend found in those works to much longer context lengths, showing that pre-training biases can be mostly overcome given enough training examples.
@@ 3 Scaling In-Context Learning
In-context learning (ICL) gives LLMs the ability to learn new tasks from examples provided only at inference time. During ICL, the LLM receives a prompt containing a set of input-output examples, also called shots, that illustrate the desired task. At the end of the prompt, we append a test input and allow the LM to make a prediction just by conditioning on the prompt and predicting the next tokens auto-regressively. Contrary to task-specific fine-tuning, ICL does not require optimizing any model parameters, allowing LLMs to perform a variety of tasks at inference.
Many-shot ICL
@@
Recent increase in context windows of LLMs allow using many more shots for ICL than typically used. Compared to few-shot ICL, many-shot learning can use larger fractions of available fine-tuning datasets. This could make task-specific fine-tuning less essential or, in some cases, even unnecessary, allowing LLMs to tackle a wider range of tasks without specialization. Furthermore, exposure to diverse ICL examples can potentially lead to better generalization, the ability to handle more complex problems than what is possible with few-shot ICL, and greater control over model outputs, potentially reducing biases stemming from pre-training.
Evaluation details
@@
We use the Gemini 1.5 Pro [^1] This corresponds to original version in the Gemini 1.5 Tech Report, released in February 2024. We note that the Gemini 1.5 Pro API now serves a newer version starting from April 2024. (Gemini Team, 2024) MoE model with 1 million token context length, the largest publicly available so far. Unless specified otherwise, we use greedy decoding for evaluation. For reliable results, we randomly sample in-context examples for each $K$-shot prompt multiple times using multiple random seeds and report average performance across seeds, along with some visualization for performance on individual seeds. To ensure that using more shots provides additional information, any $K$-shot prompt in our setup includes all in-context examples from prompts with less than $K$ examples.
3.1 Machine Translation (MT)
To evaluate how performance scales as we increase the number of in-context examples, we consider machine translation from English to a low-resource target language, where knowledge learned during pre-training may not be sufficient. Following the analysis of Robinson et al. (2023), we use the target languages with the largest gap reported between LLMs and Google Translate, namely Tamil and Kurdish, from FLORES-200 MT benchmark (NLLB Team, 2022). We modify the default 1-shot MT prompt from Gemini Team (2023) to include multiple translation pairs as shots from the FLORES dev split (containing 997 examples). We evaluate performance on the first 150 sentences from the test set using chRF, a standard MT metric based on character and word $n$-gram overlap between model-generated and reference translation. See Figure 3 for results.
Similar to Robinson et al. (2023), we observed marginal gains in the few-shot regime from 1-shot to 10-shot. However, when using the entire dev set for many-shot ICL, we observe improvements of 4.5% on Kurdish and 1.5% on Tamil, relative to the 1-shot Gemini prompt. Notably, these many-shot results establish the new-state-of-art for these language pairs, outperforming Google Translate.
@@ 3.1 Machine Translation (MT)
Figure 3: Machine Translation (MT). Test Performance improves almost monotonically as we increase the number of MT pairs provided as in-context examples during inference. In both cases, Gemini 1.5 Pro outperforms Google Translate, which obtains a chRF of 40% on Kurdish and 56% on Tamil (Robinson et al., 2023). In terms of context length, 997-shot prompt corresponds to 85.3K and 95.3K tokens for translation to Kurdish and Tamil respectively. See an example prompt in Figure A.18. These results complement the Kalamang translation with Gemini 1.5 (Gemini Team, 2024) from a single book.
3.2 Summarization
We now consider abstractive summarization, which tests the comprehension ability of LLMs to capture essence of the text. Scaling examples for ICL reveals how effectively the model can generalize this ability across different kinds of topics, a realistic scenario where adaptability is crucial. Our evaluation leverages the XSum task from the GEM benchmark (Akter et al., 2023). Using XSum development set examples (news articles and summaries) for many-shot ICL, we also investigate generalization capabilities to XLSum (Hasan et al., 2021). To assess performance, we report performance on the first 150 articles in the test set using ROUGE-L (Lin, 2004), which measures the longest common subsequence between reference and generated summaries.
As depicted in Figure 4, many-shot ICL achieves performance remarkably close to specialized summarization models fine-tuned on XSum (PEGASUS) and XLSum (mT5). Notably, XSum performance declines with more than 50 in-context examples. We surprisingly observed the many-shot prompted model occasionally generating summaries with fabricated dates and times (§A.2), despite the absence of such data in the in-context summaries. Nonetheless, performance on XLSum monotonically improves with more shots, demonstrating positive transfer from many-shot learning to a related task.
@@ 3.2 Summarization
Figure 4: Summarization. As we increase the number of (news article, summary) pairs from XSum dev set as in-context examples, XSum performance improves up to 50 shots and then deteriorates. In contrast, XLSum performance typically improves with more shots from XSum. The 500-shot prompt corresponds to 205K tokens. PEGASUS (Zhang et al., 2020) and mT5 (Hasan et al., 2021) are specialized models fine-tuned for summarization. See an example prompt in Figure A.19.
@@ 3.3 Planning: Logistics Domain
Figure 5: In-context Planning. Success rate quickly improves with up to 10 shots (37K tokens), followed by saturation up to 400 shots and a sudden performance jump at 800 shots. As a reference, we report 1-shot GPT-4 results from Valmeekam et al. (2024).
@@
While LLMs have demonstrated remarkable reasoning abilities, their ability to perform commonsense planning seems to be limited (Valmeekam et al., 2024). To this end, we verify whether many-shot ICL can improve their ability to generate simple plans on the Logistics domain, a widely used benchmark. The objective in this domain is to transport packages within cities via trucks, and between cities via airplanes. We generate a set of planning problems with 2-3 cities, 1-2 packages, 1 truck and airplane per city using a formal planning language (PDDL) generator (Seipp et al., 2022), resulting in 1.3K problems for learning and 600 for evaluation. To compute optimal solutions for each problem, we use the Fast-Downward planner (Helmert, 2006). See Figure A.20 for an example 1-shot prompt.
@@
As shown in Figure 5, in-context learning achieves significant improvement in success rate with increasing numbers of shots. While far from state-of-the-art planning approaches (e.g., Fast-Downward), our results demonstrate the potential of many-shot ICL to improve the commonsense planning abilities of LLMs.
@@
3.4 Reward Modelling: Learning Code Verifiers In-Context
A standard approach to improve LLM reasoning is to use test-time verification (Cobbe et al., 2021; Ni et al., 2023; Hosseini et al., 2024). Specifically, an LLM generates multiple candidate solutions for a given problem and a learned verifier, also known as an outcome reward model, ranks these solutions and selects the best one. Such verifiers are typically obtained by fine-tuning LLMs to predict solution correctness via binary classification. Here, we focus on learning such verifiers in-context.
To create in-context examples for code verification, we utilize correct and incorrect code-based solutions sampled from Gemini 1.0 Pro (Gemini Team, 2023) on the GSM8K train set. In the prompt, each (problem, solution) pair is appended with the question “Is the solution correct?” followed by the Yes or No token according to ground truth correctness. At inference, we modify each test (problem, solution) pair in the same way and record the logit of the Yes and No tokens (denoted by $L_{Yes}$, $L_{No}$). To compute the verifier score, we use the normalized probability of the Yes token: $\text{I\kern-1.49994ptP}(Yes)=\exp(L_{Yes})/\big(\exp(L_{Yes})+\exp(L_{No})\big)$. To evaluate verifier performance, we perform best-of-4 selection using the verifier score on 200 problems from GSM8K test set with Gemini 1.0 solutions.

### v2

@@ Abstract
Large language models (LLMs) excel at few-shot in-context learning (ICL) – learning from a few input-output examples (“shots”) provided in context at inference, without any weight updates. Newly expanded context windows allow us to investigate ICL with hundreds or thousands of examples – the many-shot regime. Going from few-shot to many-shot, we observe significant performance gains across a wide variety of generative and discriminative tasks. While promising, many-shot ICL can be bottlenecked by the available amount of human-generated outputs. To mitigate this limitation, we explore two settings: (1) “Reinforced ICL” that uses model-generated chain-of-thought rationales in place of human rationales, and (2) “Unsupervised ICL” where we remove rationales altogether, and prompt the model only with domain-specific inputs. We find that both Reinforced and Unsupervised ICL can be effective in the many-shot regime, particularly on complex reasoning tasks. Furthermore, we demonstrate that, unlike few-shot learning, many-shot learning is effective at overriding pretraining biases, can learn high-dimensional functions with numerical inputs, and performs comparably to fine-tuning. Finally, we reveal the limitations of next-token prediction loss as an indicator of ICL performance.
@@ 1 Introduction
Figure 1: Many-shot vs Few-Shot In-Context Learning (ICL) across several tasks. Many-shot ICL consistently outperforms few-shot ICL, particularly on difficult non-natural language tasks. Optimal number of shots for many-shot ICL are shown inside the bar for each task. For few-shot ICL, we either use typical number of shots used on a benchmark, for example, 4-shot for MATH, or the longest prompt among the ones we tested with less than the GPT-3 context length of 2048 tokens. Reasoning-oriented tasks, namely MATH, GSM8K, BBH, and GPQA use chain-of-thought rationales. For translation, we report performance on English to Bemba, summarization uses XLSum, MATH corresponds to the MATH500 test set, and sentiment analysis results are reported with semantically-unrelated labels. See §2, §3, and §4 for more details.
@@
A limiting factor for in-context learning (ICL) in LLMs is the context window, restricting prior research to the few-shot ICL regime. Many-shot learning – ICL with a large number of shots, for example, hundreds or thousands – allows for better task specification, can reduce the need for fine-tuning, and potentially make LLMs more versatile and adaptable. Exploring many-shot ICL is now feasible, given the recent increase in context windows of publicly available LLMs by at least $100\times$: from only a few thousand tokens in GPT-3 [8] and Llama 2 [57] to 1M tokens in Gemini 1.5 Pro [16].
@@
In this paper, we investigate how scaling the number of shots affects ICL performance on a wide variety of tasks (§2): problem solving using MATH [23] and GSM8K [10], question-answering [GPQA, 52], summarization using XSum [43] and XLSum [20], algorithmic reasoning [BBH, 56], reward modeling [Code Verifier, 24], low-resource machine translation [FLORES, 18], planning [Logistics, 54], and sentiment analysis [FP, 40]. Compared to few-shot ICL, many-shot learning performs significant better across these tasks, using several hundreds or thousands of shots (Figure 1). Furthermore, maximum performance is often achieved only once the number of shots reaches up to hundreds of thousands of tokens (Figure 2). Concurrent to our work, recent works explore many-shot ICL to jailbreak LLMs [2] (up to 256 shots) and tackle NLP classification tasks [6] (up to 80K tokens). In our work, we focus on a much wider range of tasks, use a lot more examples (up to 8192 shots), and much longer context lengths (up to 1M tokens). See §5 for a detailed discussion of related work.
@@
While many-shot ICL holds significant promise, it can be constrained by the need for high-quality, human-generated outputs. To overcome this, we introduce reinforced ICL and unsupervised ICL (§3). Inspired by the efficacy of model-generated solutions for fine-tuning [55], Reinforced ICL involves replacing human-written rationales with model-generated ones, filtered via answer correctness, for in-context learning. Inspired by task-recognition view of ICL [66], we also introduce Unsupervised ICL where we prompt the model with only problems instead of problem-solution pairs. On problem-solving tasks such as MATH, GPQA and Big-Bench Hard, we find that both reinforced and unsupervised ICL with many-shots can be more effective than few-shot ICL with human-generated rationales, with reinforced ICL being more broadly effective.
@@
Finally, we empirically study how the learning dynamics of in-context learning changes from few-shot to the many-shot regime (§4). We find that with sufficient examples, ICL can overcome pre-training biases, perform comparably to full fine-tuning, and solve high-dimensional prediction tasks with numerical inputs, namely sequential parity prediction and linear classification. This suggests the potential of many-shot ICL to adapt to unseen tasks and domains that might be misaligned with an LLM’s training data. Surprisingly, the order of examples can influence many-shot performance (§4.3) Finally, we demonstrate that long-context scaling laws [2, 68, 27] based on next-token prediction loss may not reliably predict ICL performance on problem-solving and reasoning tasks.
@@
  • Scaling ICL2): We systematically evaluate ICL performance at different scales of in-context examples for a wide range of tasks with Gemini 1.5 Pro. Our results indicate large performance jumps when transitioning from few-shot to many-shot regime.
  • Reinforced and Unsupervised ICL3): We find that using model-generated rationales or only problems can reduce the dependence of many-shot ICL on human-generated data.
  • Analysing ICL4): We show that many-shot ICL can overcome pre-training biases, perform comparably to fine-tuning, and learn non-NLP prediction tasks, where few-shot ICL struggles. We also reveal that next-token prediction loss may not be a good predictor of ICL performance.
@@ 2 Scaling In-Context Learning
During in-context learning (ICL), the LLM receives a prompt containing a set of input-output examples, also called shots, that illustrate the desired task. At the end of the prompt, we append a test input and allow the LM to make a prediction just by conditioning on the prompt and predicting the next tokens auto-regressively. Recent increase in context windows of LLMs allow using many more shots for ICL than typically used. Exposure to many more shots can lead to better generalization, handle more complex problems than what is possible with few-shot ICL, make fine-tuning less essential, and greater control over model outputs, potentially reducing biases stemming from pre-training.
Evaluation
@@
We evaluate the many-shot performance of Gemini 1.5 Pro [^1] This corresponds to original version in the Gemini 1.5 Tech Report, released in February 2024. We note that the Gemini 1.5 Pro API now serves a newer version starting from April 2024. [16] model with 1 million token context length, the largest publicly available so far. Unless specified otherwise, we use greedy decoding. For reliable results, we randomly sample in-context examples for each $K$-shot prompt multiple times using different random seeds and report average performance, along with some visualization for performance on individual seeds. To ensure that using more shots provides additional information, any $K$-shot prompt in our setup includes all in-context examples from prompts with less than $K$ examples. To reduce the inference cost, we use KV caching [49]. Next, we study many-shot ICL on typical LLM use-cases (also see §2.4).
2.1 Machine Translation
@@
Figure 3: Machine Translation (MT). Test Performance improves monotonically as we increase the number of MT pairs provided as in-context examples during inference. Notably, many-shot ICL outperforms state-of-the-art chRF2++ scores of 35% (NLLB) on Bemba and 40% (Google Translate) on Kurdish [53]. We note that 997-shot prompt corresponds to around 85K tokens. See an example prompt in Figure A.1.
@@
Figure 4: Summarization. As we increase the number of shots from XSum dev set, XSum test performance improves up to 50 shots and then deteriorates. In contrast, XLSum performance typically improves with more shots from XSum. The 500-shot prompt corresponds to 205K tokens. PEGASUS [71] and mT5 [20] are specialized models fine-tuned for summarization. See an example prompt in Figure A.2.
We consider translation from English to a low-resource target language, where many-shot ICL can complement the existing knowledge within the LLM. We use the target languages with the largest gap reported between LLMs and state-of-the-art systems [53], namely Bemba and Kurdish, from FLORES-200 benchmark [45]. We modify the default 1-shot MT prompt from Gemini Team [15] to include multiple translation pairs as shots from the FLORES dev split (containing 997 examples). We evaluate performance on the first 150 sentences from the test set using chrF2++ [50], a standard metric based on character and word $n$-gram overlap between generated and reference translation.
See Figure 4 for results. Similar to Robinson et al. [53], we observed small gains in the few-shot regime from 1-shot to 10-shot, particularly on Kurdish. However, when using the entire dev set for many-shot ICL, we observe improvements of 15.3% on Bemba and 4.5% on Kurdish, relative to the 1-shot Gemini prompt. Overall, these results establish the new-state-of-art for these language pairs.
2.2 Abstractive Summarization
To investigate how scaling ICL examples can impact the comprehension ability of LLMs, we now consider abstractive news summarization using XSum dataset from the GEM benchmark [1]. Using XSum dev set examples containing news articles and summaries, we also evaluate how many-shot ICL generalizes to XLSum [20]. We report performance on 150 test articles using ROUGE-L [35], which measures the longest common subsequence between reference and generated summaries.
As depicted in Figure 4, peak performance with many-shot ICL is remarkably close to specialized models fine-tuned on XSum and XLSum. However, XSum performance declines with more than 50 in-context examples. Surprisingly, we observed the many-shot prompted model occasionally generating summaries with fabricated dates and times (§A.8), despite the absence of such data in the in-context summaries. Nonetheless, performance on XLSum monotonically improves with more shots, demonstrating positive transfer from many-shot learning to a related task.
@@ 2.3 Planning: Logistics Domain
Figure 5: In-context Planning. Success rate quickly improves with up to 10 shots (37K tokens), followed by saturation up to 400 shots and a sudden performance jump at 800 shots. As a reference, we report 1-shot GPT-4 results from Valmeekam et al. [59]. See Figure A.3 for an example 1-shot prompt.
@@
Recent work has highlighted shortcomings in planning abilities of LLMs [59]. To this end, we evaluate whether many-shot ICL can improve their ability to generate simple plans on the Logistics domain, a widely used benchmark. The objective in this domain is to transport packages within cities via trucks, and between cities via airplanes. We generate a set of planning problems with 2-3 cities, 1-2 packages, 1 truck and airplane per city using a formal planning language (PDDL) generator [54], resulting in 1.3K problems for learning and 600 for evaluation. To compute optimal solutions for each problem, we use the Fast-Downward planner [21].
@@
As shown in Figure 5, we observe significant improvement in success rate with increasing numbers of ICL shots. While far from state-of-the-art planning approaches (e.g., Fast-Downward), our results demonstrate the potential of many-shot ICL to improve the commonsense planning abilities of LLMs.
@@
2.4 Reward Modelling with Many-Shot ICL: Learning Code Verifiers